OpenStudy (anonymous):

PLEASE HELP ME ! Let f(x) = 3x+2 and g(x) = x2 Part 1 [2 points] Find f(g(x)) Part 2 [2 points] Find g(f(x)) Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2.

OpenStudy (anonymous):

f(g(x): Plug in g(x) into f(x) [the value of g(x) is x in f] g(f(x): Plug in f(x) into g(x) [the value of f(x) is x in g] f(x) in words: f is a function of x (f is dependent on x) g(x) in words; g is a function of x f(g(x) in words: f is a function of g g(f(x) in words: g is a function of f

OpenStudy (anonymous):

how do I do that ?

OpenStudy (anonymous):

sorry, I suck at these

OpenStudy (anonymous):

f(x) = 3x+2 and g(x) = x2 fgx I'd do the first question just for example: f(x) = 3x + 2 Plug in g(x) as x: f(gx) = 3(x2) + 2

OpenStudy (anonymous):

can you pleaseee help me with all of these

OpenStudy (anonymous):

@dauspex

OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x2 Part 1 [2 points] Find f(g(x)) f(x) = 3x + 2 f(gx) = 3(x2) + 2 = 3x2 +2 Part 2 [2 points] Find g(f(x)) g(x) = x2 g(fx) = (3x+2)^2 Part 3 [4 points] Use complete sentences to explain any difference in part 1 and part 2. f(g(x) in words: f is a function of g g(f(x) in words: g is a function of f

OpenStudy (anonymous):

thank you soo much !